OpenStudy (anonymous):

One end of a rope is fastened to a boat and the other end is wound around a windlass located on a dock at a point 4m above the level of the boat. If the boat is drifting away from the dock at the rate of 2m/min, how fast is the rope unwinding at the instant when the length of the rope is 5m?

OpenStudy (anonymous):

looks like pythagoras to me

OpenStudy (anonymous):

yup. i think its pythagoras.. but some how.. i was confused with the question.. i cant even understand how what it want and how to start.. ><

OpenStudy (anonymous):

if we call the distance from the boat to the dock x, and the distance from the boat to the windlass (whatever that is) y, then we have \[4^2+x^2=y^2\]

OpenStudy (anonymous):

draw a picture and you will have a triangle with hypotenuse the length of the rope (i called that y) one leg 4 and the other the distance from the dock to the boat, (i called that x)

OpenStudy (anonymous):

in my use of variables you are told \[x'=2\] because x is the distance from the boat to the dock. and you are asked for \[y'\] the speed at which the ropes is increasing

OpenStudy (anonymous):

starting with the equation \[4^2+x^2=y^2\] and differentiating wrt time we get \[2xx'=2yy'\] solving for y' gives \[y'=\frac{xx'}{y}\]

OpenStudy (anonymous):

now plug in the numbers. we are given ."If the boat is drifting away from the dock at the rate of 2m/min" meaning \[x'=2\] and also "how fast is the rope unwinding at the instant when the length of the rope is 5m " tells us \[y=5\]

OpenStudy (anonymous):

pythagoras tells us that if y = 5 then x = 3 (we have a nice 3-4-5 right triangle) and so \[y'=\frac{3\times 2}{5}=\frac{6}{5}\]

OpenStudy (anonymous):

meters per minute or whatever the units are

OpenStudy (anonymous):

steps more or less clear?

OpenStudy (anonymous):

Clear!! thx!

OpenStudy (anonymous):

yw